### v1

@@ 3 Supernova remnants and pulsar wind nebulae
The young pulsar PSR B0540-69 is the first to display an increase in brightness in its PWNe following a change in its spin-down rate [77]. This discovery offers valuable insights into the correlation between the central pulsar engine and the fluctuations of the nebula brightness. Consequently, understanding the relationship between young pulsars and their surrounding PWNe is essential for examining the braking mechanisms of pulsars, as discussed by [77]. If the enhanced energy is converted to more particles, these particles could be injected into the surrounding environment and affect the original structure of PWNe. This process might change the polarimetry feature of PWNe gradually, which could be observed and studied utilizing the observations of eXTP-PFA. X-ray polarization detection could help us to understand the temporal evolution and formation of the PWNe structure affected by the enhanced particle. IXPE has detected a PD of 24.5% 5.3% at 4-6 keV in the off-pulse phase bin, which covers 35% of the entire phase [74]. Using the same exposure of 850 ks, eXTP would provide a 15 detection with a smaller PD error of 1.6%. Assuming that the fluence increases by 30% and the increased portion is not polarized, the PD X-ray would decrease 5%. In this case, the decrease is difficult to detect, requiring 7 Ms to achieve a 3 detection (the significance of the polarization variant). If the outburst disturbs the magnetic fields, it may lead to a larger drop in the PD. Taking into account a decrease of 10%, we could achieve a 3 detection with an exposure of 195 ks. For a 15% decrease, a 5 detection would be achieved with only 260 ks observation. The same could be applied to other PWNe with the central pulsar, which shows flux increase or outburst behavior. For example, the high magnetic field pulsar PSR J18460258 in the young SNR Kes 75 exhibited a magnetar-like outburst [78] and the pulsar flux increased by 6 times with an emerging thermal component [79]. Rapid follow-up of similar events with eXTP could reveal the corresponding changes in the PWN magnetic field.
@@
Supernovae regulate the chemical evolution of galaxies via the ejection of heavy elements into the interstellar medium. However, the total chemical output of supernovae remains highly uncertain. These uncertainties come from ongoing debates regarding which stellar systems ultimately result in successful supernova explosions and the identification of viable mechanisms responsible for the production of SNe. Metal measurements of supernovae and their remnants are thus of great importance in addressing the above debates, as they provide direct observational constraints. The metal yields of an SN depend on the progenitor and explosion mechanism, and thus the metal measurements provide a viable way to find out which systems can produce supernovae [80].
Metal measurements of SNRs are mostly conducted using X-ray spectroscopic observations, which are proven to be an efficient way to constrain the physical and chemical parameters of the SN materials. SFA of eXTP works in the 0.5–10 keV, covering the lines of elements from Nitrogen to Nickel. By comparing the measured abundances with those predicted by nucleosynthesis models, we can build an SNR-progenitor connection.
@@
The left panel of Figure 5 shows a simulated spectrum of SNR W49B with only 5 ks SFA observation, based on a model obtained by fitting 270 ks Chandra spectra [85]. W49B has been extensively studied in X-rays, but its origin remains a subject of debate, with ongoing discussions on whether it is a thermonuclear or a core-collapse SNR. With 5 ks observation, SFA can clearly detect faint Cr and Mn lines, Fe lines at different ionization states, and the radiative recombination continuum (RRC) of Fe. By fitting the spectrum with a recombining plasma model , we found that the metal abundances of Cr, Mn, and Fe can be constrained with a precision of ( for Ni, 1-). This verifies the ability of the SFA to measure the metal composition and physical properties of SNRs. Note that SFA-Ts is not designed for spatially resolved studies, given its large pixel size, so the spectrum is simulated for the whole SNR. SFA-I can provide an angular resolution of , which will greatly benefit the observations of SNRs by resolving structures and detecting faint features.
eXTP can establish an SNR-progenitor connection for a sample of SNRs and provide valuable insight into the SN progenitor and chemical outputs. Its large effective area and moderate FOV () are suitable for observing small and medium-sized SNRs, although mosaic observations are also possible for some large SNRs. Some of these SNRs exhibit elevated metal abundances as revealed with XMM-Newton and/or Chandra observations, but their spectra, particularly in the hard X-ray regime, suffer from insufficient statistical quality. SFA will be able to complete a survey of these SNRs and provide spectra with high statistical quality within a manageable time. For example, the right panel of Figure 5 shows the simulated spectrum of SNR G306.3 with a 50-ks SFA observation, based on an absorbed model described by [86]. This SNR is proposed as a remnant of a Ca-rich transient, which is a subgroup of peculiar thermonuclear SNe. Our simulation shows that the abundance of Ca and Fe will be determined with an accuracy of 20% (90% confidence level), and other elemental abundances will have an accuracy better than 10%. This precision will help distinguish the explosion mechanisms and properties of the progenitor system.
@@
What is the gamma-ray production mechanism of the SNR-MC system? Is it produced by direct or indirect collision? Several dozen SNR-MC systems have been detected so far in the gamma-ray band, and their radiation is generally believed to be produced by hadrons accelerated by shock waves bombarding molecular clouds (MCs) [e.g., 87, 88]. However, the bombardment process remains a topic of debate. There are currently two scenarios. One is that the particles accelerated by the shock wave are confined near the shock. As the shock wave expands and reaches the MC, energetic hadrons collide with the MCs to produce hadronic gamma rays, namely direct collision [e.g., 89, 90, 91]. The other is that the accelerated particles escape the acceleration site, and impact the MCs before the shock wave arrives. Gamma rays come from the accumulation of hadrons that escape from the shock wave, that is, indirect collision or escaping collision[e.g., 92, 93, 94, 95, 96, 97]. Both types of models can explain some of the observed gamma-ray spectra, but the predicted spatial distributions are different. The gamma rays of the direct collision model are mainly concentrated near the shock, while the gamma rays of the escape model can be in the far upstream area of the shock. Due to the limited angular resolution of gamma-ray telescopes ( degree), it is difficult to distinguish the two models with current gamma-ray observations. Due to the better spatial resolution in the X-ray band, it is possible to distinguish the two models from the spatial distribution by searching for synchrotron X-rays of secondary particles from hadronic processes. For an MC with a mass of 10,000 and a distance of 30 pc from the center of the SNR, the synchrotron X-ray flux of secondary particles from hadronic processes can reach a few erg if the SNR is within a few thousand years old and within a few kpc from the Earth. This emission may be detected by the SFA because of its large collection area.
With the discovery of ultra-high-energy (UHE) gamma-ray sources by the LHAASO and HAWC experiments, determining whether this radiation is leptonic or hadronic still requires further multi-wavelength observations. If this gamma-ray emission originates from the decay of mesons produced in SNR-MC systems, secondary electrons can potentially emit synchrotron photons in the X-ray band [98], which could be tested by the eXTP SFAs, as well as provide a strong constraint on the primary CR spectrum.
@@ 4 Cataclysmic variables
eXTP can address the above issues by studying selected samples of CVs. The SFA will be capable of detecting low-amplitude fast periodic and aperiodic variabilities in both magnetic and nonmagnetic systems, as well as performing time-resolved spectroscopy. The PFA will be able to detect potential X-ray polarization signatures in bright magnetic CVs during their high states.
How does mass ejection occur in nova explosions?
@@
Novae are, after thermonuclear supernovae, the second most luminous X-ray sources among accreting WDs, with emission spanning from very soft to hard X-rays. The soft X-rays probe continuous nuclear burning on the WD surface, while the hard X-rays are thought to trace the initial shocks within the ejecta and between the ejecta and the surrounding circumstellar material, as well as the onset of resumed accretion. However, the details of the mass outflow and the shaping and evolution of the ejected material are poorly understood. Detection of high-energy gamma rays ( MeV) from an increasing number of novae by the Fermi LAT [120, 121] suggests that these systems are sites of particle acceleration. This was previously predicted for RS Oph [122], a recurrent symbiotic nova, where strong shocks between the nova ejecta and the red giant wind were detected in hard X-rays by RXTE [123]. The recent detection of this nova at even much higher energies up to TeV [124, 125] indicates that symbiotic systems are crucial testbeds for studying particle acceleration. In classical novae, however, there is no red giant wind interacting with the nova mass outflow, as the companion star is a main-sequence star. Therefore, particle acceleration in these systems is more difficult to explain [126]. Prompt observations of nova explosions by eXTP would be crucial for a better understanding of these processes in both types of novas (see Fig. 6).
@@
It is crucial to observe the spectral evolution of nova explosions across the broadest X-ray range, from the initial hard X-ray onset to the later super-soft X-ray phase [123, 127, 128]. Furthermore, X-ray variability and periodicities on a variety of timescales, from days to sub-minute oscillations, remain poorly understood. Short oscillations in soft X-rays may originate from H burning instabilities [113], and the short-time variability in the harder component of the ejecta and/or the accretion disc is still to be investigated. The wide-field W2C with an exposure time of one or several days is expected to detect the early X-ray emissions from bright novae, such as the recurrent nova RS Oph, which has a hard X-ray flux (20 – 100 keV) of 10-10 erg cm-2 s-1 during the initial days of its outburst [129]. The SFA will then track the X-ray spectral evolution from the initial hard X-ray phase to the later super-soft X-ray phase and investigate the characteristics of the fast variability in these bright novae. The large effective area of SFAs with a wonderful time resolution, and the high orbit of eXTP combine nicely for exploring the variability of novae in outburst and in particular the short-time oscillations.
@@
Dwarf nova (DN) outbursts are thought to be the result of disk instabilities, but there is still limited understanding of the changes that occur at the inner boundary layer. The boundary layer is optically thin during quiescence and becomes optically thick during outburst [130]. A suppression of hard X-rays during optical outbursts was previously believed to be a universal feature. However, the few DNs observed so far have shown considerable diversity, and not all exhibit the expected X-ray/optical anti-correlation [131]. Understanding the fundamental parameters that distinguish different regimes of the boundary layer is essential. Another challenge is the detection of radio emission from DNs and CVs of high mass accretion rate [132, 133, 134], which has been interpreted as evidence of a radio jet. CVs were once thought to be incapable of launching jets, but recent studies have shown that nova-like CVs can have optically thin and inefficient accretion disk boundary layers [135], allowing the retention of enough energy to power jets. If this is the case, a radio/X-ray correlation, similar to those observed in neutron star and black hole binaries [136], should be present in these systems.
eXTP can tackle these complex issues by monitoring a select few DNe with comprehensive multi-outburst coverage, using time-resolved spectroscopy to explore these systems.
@@
Period bouncers are CVs that have evolved past the orbital period minimum (80 min), where the donor star becomes degenerate (brown dwarf-like) and the system “bounces back” to longer orbital periods. Theoretical models predict that 40–80% of CVs should be period bouncers [e.g., 137]. Their cumulative X-ray emission could significantly contribute to the Galactic Ridge X-ray Emission (GRXE), but their low individual luminosities ( erg s-1) require deep surveys for detection [138, 139]. However, observations reveal that only 7–14% of the observed CVs are period bouncers [140, 141], creating a significant discrepancy. The main challenges include their intrinsic faintness and low mass transfer rates ([141]. Recent works have proposed more than 100 candidates, several of which have been confirmed by X-ray surveys (e.g., eROSITA) by detecting accretion signatures [142]. However, whether the majority of the candidates are truly period bouncers remains an open question because of the lack of X-ray detection and/or X-ray spectroscopy.
The confirmation of a period bouncer depends on the detection of X-ray emission, which signals ongoing accretion. Although H emission is commonly used to identify CVs, this method is unreliable for period bouncers due to contamination of chromospheric activity in their low-mass donors [142]. A dedicated survey on candidate period bouncers with eXTP would be crucial for probing their properties and population. With adequate X-ray exposure (e.g. 50 ks exposure time for sources with a distance of 500 pc), spectroscopic and timing analyses will provide critical constraints on period bouncer physics, particularly their accretion mechanisms and WD properties. X-ray spectra can constrain the maximum emission temperature, while timing analysis reveals the accretion geometry (e.g., disk-fed for nonmagnetic WDs versus column accretion in magnetic systems). The WD mass can then be derived via the - relation [143]. Such a survey would also constrain magnetic field strengths, key parameters for testing binary evolution models, particularly the crystallization-driven dynamo scenario [144]. Considering SFA’s large effective area and its good sensitivity to point sources, eXTP is capable of detecting X-ray emission of candidate period bouncers within a distance range of 500-1000 parsecs.
@@ 5.1 Low-mass X-ray binaries
Low-mass X-ray binaries (LMXBs) are interacting binary systems in which a compact object like a neutron star or a black hole, accretes matter from a low-mass companion star (typically ). The accreted material, transferred via Roche lobe overflow, forms an accretion disk that radiates strongly in X-rays as gravitational potential energy is converted into heat near the compact object. These systems exhibit a broad range of temporal and spectral phenomena, including QPOs, relativistic jets, thermonuclear bursts & pulsations in neutron star accretors. LMXBs serve as laboratories for studying accretion physics, strong gravity and magnetic fields, and the evolution of compact binaries, and they include diverse sub-classes such as black hole X-ray binaries (BHXRBs), accreting millisecond X-ray pulsars (AMXPs), transitional millisecond pulsars (tMSPs), and nuclear–powered bursting neutron stars.

### v2

@@ 3 Supernova remnants and pulsar wind nebulae
The young pulsar PSR B0540-69 is the first to display an increase in brightness in its PWN following a change in its spin-down rate [77]. This discovery offers valuable insights into the correlation between the central pulsar engine and the fluctuations of the nebula brightness. Consequently, understanding the relationship between young pulsars and their surrounding PWNe is essential for examining the braking mechanisms of pulsars, as discussed by [77]. If the enhanced energy is converted to more particles, these particles could be injected into the surrounding environment and potentially affect the structure of PWNe. This process might gradually change the polarimetry properties of PWNe, which could be observed and studied utilizing the observations of eXTP-PFA. X-ray polarization measurements could help us to understand the temporal evolution and formation of the PWNe structure affected by the enhanced particle flow. IXPE has detected a PD of 24.5% 5.3% at 4-6 keV in the off-pulse phase bin, which covers 35% of the entire phase [74]. Using the same exposure of 850 ks, eXTP would provide a 15 detection with a smaller PD error of 1.6%. Assuming that the fluence increases by 30% and the increased portion is not polarized, the PD X-ray would decrease 5%. In this case, the decrease is difficult to detect, requiring 7 Ms to achieve a 3 detection. If the outburst disturbs the magnetic fields, it may lead to a larger drop in the PD. Assuming a decrease of 10%, we could achieve a 3 detection with an exposure of 195 ks. For a 15% decrease, a 5 detection would be achieved with only 260 ks observation. The same could be applied to other PWNe the central pulsars, which show flux increase or outburst behavior. For example, the high magnetic field pulsar PSR J18460258 in the young SNR Kes 75 exhibited a magnetar-like outburst [78] and the pulsar flux increased by 6 times with an emerging thermal component [79]. Rapid follow-up of similar events with eXTP could reveal the corresponding changes in the PWN magnetic field.
@@
Supernovae regulate the chemical evolution of galaxies via the ejection of heavy elements into the interstellar medium [80]. However, the total chemical output of supernovae remains highly uncertain. These uncertainties come from ongoing debates regarding which stellar systems ultimately result in successful supernova explosions and the identification of viable explosion mechanisms. Metal measurements of supernovae and their remnants are thus of great importance in addressing the above debates, as they provide direct observational constraints. The metal yields of an SN depend on the progenitor and explosion mechanism, and thus the metal measurements provide a viable way to constrain which systems can produce supernovae [80].
Metal measurements of SNRs are mostly conducted using X-ray spectroscopic observations, which are proven to be an efficient way to constrain the physical and chemical parameters of the SN materials. SFA of eXTP works in the 0.5–10 keV range, covering the lines of elements from Nitrogen to Nickel. By comparing the measured abundances with those predicted by nucleosynthesis models, we can build an SNR-progenitor connection.
@@
The left panel of Figure 5 shows a simulated spectrum of SNR W49B with only 5 ks SFA observation, based on a model obtained by fitting 270 ks Chandra spectra [85]. W49B has been extensively studied in X-rays, but its origin remains a subject of debate over whether it is a thermonuclear or a core-collapse SNR. With 5 ks observation, SFA can clearly detect faint Cr and Mn lines, Fe lines at different ionization states, and the radiative recombination continuum (RRC) of Fe. By fitting the spectrum with a recombining plasma model , we found that the metal abundances of Cr, Mn, and Fe can be constrained with a precision of ( for Ni, 1-). This verifies the ability of the SFA to measure the metal composition and physical properties of SNRs. Note that SFA-Ts is not designed for spatially resolved studies, given its large pixel size, so the spectrum is simulated for the whole SNR. SFA-I can provide an angular resolution of , which will greatly benefit the observations of SNRs by resolving structures and detecting faint features.
eXTP can establish an SNR-progenitor connection for a sample of SNRs in our Galaxy and Magellanic Clouds, and provide valuable insight into the SN progenitor and chemical outputs. Its large effective area and moderate FOV () makes it well-suited for observing small and medium-sized SNRs, although mosaic observations are also possible for some large SNRs. Some of these SNRs exhibit elevated metal abundances as revealed with XMM-Newton and/or Chandra observations, but their spectra, particularly in the hard X-ray regime, suffer from insufficient statistical quality. SFA will be able to conduct a survey of these SNRs and provide spectra with high statistical quality within a manageable time. For example, the right panel of Figure 5 shows the simulated spectrum of SNR G306.3 with a 50-ks SFA observation, based on an absorbed model described by [86]. This SNR is proposed as a remnant of a Ca-rich transient, which is a subgroup of peculiar thermonuclear SNe. Our simulation shows that the abundance of Ca and Fe will be determined with an accuracy of 20% (90% confidence level), and other elemental abundances will have an accuracy better than 10%. This precision will help distinguish the explosion mechanisms and properties of the progenitor system.
@@
What is the gamma-ray production mechanism of the SNR-MC system? Is it produced by direct or indirect interaction? Several dozen SNR-MC systems have been detected so far in the gamma-ray band, and their radiation is generally believed to be produced by hadrons accelerated by shock waves bombarding molecular clouds (MCs) [e.g., 87, 88]. However, the bombardment process remains a topic of debate. There are currently two scenarios. One is that the particles accelerated by the shock wave are confined near the shock. As the shock wave expands and reaches the MC, energetic hadrons collide with the MCs to produce hadronic gamma rays, namely direct collision [e.g., 89, 90, 91]. The other is that the accelerated particles escape the acceleration site, and impact the MCs before the shock wave arrives. Gamma rays come from the accumulation of hadrons that escape from the shock wave, that is, indirect collision or escaping collision[e.g., 92, 93, 94, 95, 96, 97]. Both types of models can explain some of the observed gamma-ray spectra, but the predicted spatial distributions are different. The gamma rays of the direct collision model are mainly concentrated near the shock, while the gamma rays of the escape model can be in the far upstream area of the shock. Due to the poor angular resolution of current gamma-ray telescopes ( degree), it is difficult to distinguish the two models. X-ray telescopes often have much better angular resolution and may distinguish the two models by searching for synchrotron X-rays of secondary particles produced from hadronic processes. For an MC with a mass of 10,000 and a distance of 30 pc from the center of the SNR, the synchrotron X-ray flux of secondary particles from hadronic processes can reach a few erg if the SNR is within a few thousand years old and within a few kpc from the Earth. This emission may be detected by the SFA because of its large collection area.
With the discovery of ultra-high-energy (UHE) gamma-ray sources by LHAASO, determining whether this radiation is leptonic or hadronic still requires further multi-wavelength observations. If this gamma-ray emission originates from the decay of mesons produced in SNR-MC systems, secondary electrons can potentially emit synchrotron photons in the X-ray band [98]. Detecting this emission with the eXTP SFA would test this scenario and provide strong constraints on the primary CR spectrum.
@@ 4 Cataclysmic variables
eXTP can address the above issues by studying selected samples of CVs. The SFA will be capable of detecting low-amplitude fast periodic and aperiodic variabilities in both magnetic and nonmagnetic systems, as well as performing time-resolved spectroscopy. For example, an exposure time of 30 ks or longer using the SFA enables the detection and analysis of rapid oscillations in nearby outbursting DNe (e.g., VW Hyi, U Gem, and OY Car). The PFA will be able to detect potential X-ray polarization signatures in bright magnetic CVs during their high states. For instance, the 2–10 keV flux of the magnetic CV AM Herculis in its high state is about 10-10 erg cm-2 s-1. An exposure time of 60 ksec with the PFA is expected to provide an MDP99 of for this source.
@@
Novae are, after thermonuclear supernovae, the second most luminous X-ray sources among accreting WDs, with emission spanning from very soft to hard X-rays. The soft X-rays probe continuous nuclear burning on the WD surface, while the hard X-rays are thought to trace the initial shocks within the ejecta and between the ejecta and the surrounding circumstellar material, as well as the onset of resumed accretion. However, the details of the mass outflow and the shaping and evolution of the ejected material are poorly understood. Detection of high-energy gamma rays ( MeV) from an increasing number of novae by the Fermi-LAT [120, 121] suggests that these systems are sites of particle acceleration. This was previously predicted for RS Oph [122], a recurrent symbiotic nova, where strong shocks between the nova ejecta and the red giant wind were detected in hard X-rays by RXTE [123]. The recent detection of this nova at even much higher energies up to TeV [124, 125] indicates that symbiotic systems are crucial testbeds for studying particle acceleration. In classical novae, however, there is no red giant wind interacting with the nova mass outflow, as the companion star is a main-sequence star. Therefore, particle acceleration in these systems is more difficult to explain [126]. Prompt observations of nova explosions by eXTP would be crucial for a better understanding of these processes in both types of novae (see Fig. 6).
@@
It is crucial to observe the spectral evolution of nova explosions across the broadest X-ray range, from the initial hard X-ray onset to the later super-soft X-ray phase [123, 127, 128]. Furthermore, X-ray variability and periodicities on a variety of timescales, from days to sub-minute oscillations, remain poorly understood. Short oscillations in soft X-rays may originate from H burning instabilities [113], and the short-time variability in the harder component of the ejecta and/or the accretion disc is still to be investigated. The wide-field W2C with an exposure time of one or several days is expected to detect the early X-ray emissions from bright novae, such as the recurrent nova RS Oph, which has a hard X-ray flux (20 – 100 keV) of 10-10 erg cm-2 s-1 during the initial days of its outburst [129]. The SFA will then track the X-ray spectral evolution from the initial hard X-ray phase to the later super-soft X-ray phase and investigate the characteristics of the fast variability in these bright novae. The large effective area and superb time resolution of the SFA, combined with the high orbit of eXTP, make it well-suited for exploring the variability of novae in outburst and in particular the short-time oscillations.
@@
Dwarf nova (DN) outbursts are thought to be the result of disk instabilities, but the changes that occur at the inner boundary layer are not yet fully understood. The boundary layer is optically thin during quiescence and becomes optically thick during outburst [130]. A suppression of hard X-rays during optical outbursts was previously believed to be a universal feature. However, the few DNs observed so far have shown considerable diversity, and not all exhibit the expected X-ray/optical anti-correlation [131]. Understanding the fundamental parameters that distinguish different regimes of the boundary layer is essential. Another challenge is the detection of radio emission from DNs and CVs of high mass accretion rate [132, 133, 134], which has been interpreted as evidence of a radio jet. CVs were once thought to be incapable of launching jets, but recent studies have shown that nova-like CVs can have optically thin and inefficient accretion disk boundary layers [135], allowing the retention of enough energy to power jets. If this is the case, a radio/X-ray correlation, similar to those observed in neutron star and black hole binaries [136], should be present in these systems.
eXTP can tackle these complex issues by monitoring a selected sample of DNe with comprehensive multi-outburst coverage, using time-resolved spectroscopy to explore these systems.
@@
Period bouncers are CVs that have evolved past the orbital period minimum (80 min), where the donor star becomes degenerate (brown dwarf-like) and the system “bounces back” to longer orbital periods. Theoretical models predict that 40–80% of CVs should be period bouncers [e.g., 137]. Their cumulative X-ray emission could significantly contribute to the Galactic Ridge X-ray Emission (GRXE), but their low individual luminosities ( erg s-1) require deep surveys for detection [138, 139]. However, observations reveal that only 7–14% of the observed CVs are period bouncers [140, 141], creating a significant discrepancy. The main challenges include their intrinsic faintness and low mass transfer rates [141]. Recent works have proposed approximately 200 candidates, with over 80 located within 500 pc [142]. Among these, several candidates have been confirmed by X-ray surveys (e.g., eROSITA) by detecting accretion signatures [143]. However, whether the majority of the candidates are truly period bouncers remains an open question because of the lack of X-ray detection and/or X-ray spectroscopy.
The confirmation of a period bouncer depends on the detection of X-ray emission, which signals ongoing accretion. Although H emission is commonly used to identify CVs, this method is unreliable for period bouncers due to contamination of chromospheric activity in their low-mass donors [143]. Dedicated pointing observations on strong period-bounce candidates with eXTP would be crucial for probing their properties and population. With adequate X-ray exposure (e.g. 50 ks exposure time for sources with a distance of 500 pc), spectroscopic and timing analyses will provide critical constraints on period bouncer physics, particularly their accretion mechanisms and WD properties. X-ray spectra can constrain the maximum emission temperature, while timing analysis reveals the accretion geometry (e.g., disk-fed for nonmagnetic WDs versus column accretion in magnetic systems). The WD mass can then be derived via the - relation [144]. These observations would also constrain magnetic field strengths, key parameters for testing binary evolution models, particularly the crystallization-driven dynamo scenario [145]. Considering SFA’s large effective area and its good sensitivity to point sources, eXTP is capable of detecting X-ray emission of candidate period bouncers within a distance range of 500-1000 parsecs.
@@ 5.1 Low-mass X-ray binaries
Low-mass X-ray binaries (LMXBs) are interacting binary systems in which a compact object like a neutron star or a black hole, accretes matter from a low-mass companion star (typically ). The accreted material, transferred via Roche lobe overflow, forms an accretion disk that radiates strongly in X-rays as gravitational potential energy is converted into heat near the compact object. These systems exhibit a broad range of temporal and spectral phenomena, including QPOs, relativistic jets, thermonuclear bursts, and pulsations in neutron star accretors. LMXBs serve as laboratories for studying accretion physics, strong gravity and magnetic fields, and the evolution of compact binaries, and they include diverse sub-classes such as black hole X-ray binaries (BHXRBs), accreting millisecond X-ray pulsars (AMXPs), transitional millisecond pulsars (tMSPs), and nuclear–powered bursting neutron stars.